Mopra current issues
- TCS version current should be used.
- Please note that IRC to coordinate your observing handovers is no longer used. Please use the chat window in the top left hand side of the TCS screen.
- procedural changes for generator operation
- new online observer tools
- information on how to use IRC for remote observing coordination
- guide on debogging yourself with VNC problems.
- TOAD is intermittently complaining about the Dataset RPC Server being unavailable. This can be ignored, but please notify Balt if you encounter that message.
- The conversion power supply warning in TOAD can be safely ignored until further notice.
- A new tool for our mapping observers has become available: The Beam Position Mapper. It plots RA/Dec map of all the beam positions in the given RPF files. Use this tool when mapping, as time and maps have been lost in the past due to faulty antenna driving. In these plots, such a fault is immediately obvious. Contact Balt if you need an introduction. Caution: Be careful not to use it unless you are the observer in charge. It is NOT multi user capable and plotting requests lodged while another request is in progress will be lost.
- OTF Schedule Maker version 6 has finally been released: It incorporates a graphical zoom window tool (similar to mops_vis) and automatically optimizes your center frequencies.
- TCS version next is in use currently.
Dec 2014: New Fault Report systemThe old FR system has been discontinued (with all old FRs imported into the new system for reference). Use the new link to submit a FR.
MAR 2014: New observing servers and 12mm issuesThe VNC sessions have changed and are now running on bigrock.atnf.csiro.au instead of mpvnc1. Please update your links. The 12mm system remains uncalibrated for now due to difficulties getting the RF powers under control.
SEP 2013: The Mopra Southern Galactic Plane CO SurveyThe first public data release from the Mopra Southern Galactic Plane Carbon Monoxide Survey is now available through the ATNF online data archive at atoa.atnf.csiro.au (note that an ATNF OPAL account is required for access). The Mopra CO survey is being conducted between l=305-345°, b=+/-0.5° of the principal isotopologues of the CO J=1-0 line, in particular 12CO 115 GHz and 13CO 110 GHz. The survey has 0.6 arcminute and 0.1 km/s resolution, representing an order of magnitude increase over previous CO surveys of this region. It provides the highest-resolution, wide-field view of the molecular medium of the fourth quadrant of the Galaxy now available. The first data release contains the data cubes from the first degree of the survey undertaken, of the G323 region. Further information is available from the project website and also in a recently published paper in PASA (Burton et al, 2013, "The Mopra Southern Galactic Plane CO Survey", PASA, v30, e044 doi:10.1017/pasa.2013.22). The survey itself is ongoing, with data from about 25 sq. deg of the Galactic plane obtained to date (l=317°-342°). The data cubes from this work will be made publicly available through subsequent data releases, as they are analysed and published. However, if you are interested in early access to the data set for a particular science project, please contact the project PI, Michael Burton, at m.burton [at] unsw.edu.au.
JUN 2013: Mopra characterisationA paper entitled "Characterization of the MALT90 Survey and the Mopra Telescope at 90 GHz" (arXiv:1306.0560) by Foster et al. gives a nice presentation on the performance of Mopra in the 3mm band.
MAY 2013: Mopra operational againAs outlined in this News Item the Mopra telescope, which had been out of action since bushfires burned through much of the Warrumbungle National Park on January 13th, 2013, is operational again. Fortunately, the telescope itself and the key data-taking equipment were spared by the fire, but mains power was cut and the on-site building suffered significant damage. Among the work carried out over the three months following the fire was the removal of rubble and debris from the site, replacement of electrical switchboards damaged in the fire, professional cleaning of electrical and electronic equipment to remove any particulate matter deposited during the fire, confirmation of the structural integrity of the equipment room, re-establishment of the required air-conditioning to the equipment room, and a systematic program of progressive checks of the state of the computers and equipment. This work was completed at the beginning of May, and a series of observations was then conducted to confirm the telescope was fully operational. Work to permanently repair or replace the building and other infrastructure will continue over the next month or two.
The 2013APR schedule was released on May 23rd, with block allocations
to the NAOJ and UNSW/Adelaide consortia, and allocations of ATNF time
for single dish observing and LBA participation.
AUG 2012: Mopra future operating modelA Call for expressions of interest to operate or fund the Mopra telescope from October 2012 was released on 18 November 2011. In August 2012, Phil Diamond (CASS Chief and ATNF Director) announced: "We are pleased to announce that funding has been secured to operate the Mopra Telescope for three years, commencing in October 2012. Mopra operations will initially be supported by the National Astronomical Observatory of Japan, the University of New South Wales, and the University of Adelaide. The model for allocating observing time is still under development. We will advise the community once it is finalised."
MAR 2012 / edw126: Livedata/Gridzilla webpage movedThe Livedata/Gridzilla webpage has moved to a new location.
OCT 2011 / edw126: New version of ASAPVersion 4.0 of ASAP (the ATNF Spectral Analysis Package) was released in October 2011. See also the ASAP wiki for more details.
27 AUG 2011 / ind006: TCS version and Observing coordination
30 MAY 2011 / ind006: New QRH releasedNew version of the QRH released.
01 JUN 2010 / ind006: TCS versionTCS version current should be used.
06 MAY 2010 / edw126: New version of ASAPVersion 3.0 of ASAP (the ATNF Spectral Analysis Package) was released on 06 May 2010. See also the ASAP wiki for more details.
03 FEB 2010 / ind006: New QRH releasedExtensive update to the QRH: Please read the changes, they mostly involve
01 FEB 2010 / ind006: TOAD 1.0.0 released!We're celebrating a milestone today with the release of TOAD 1.0.0! It's been in the making for well over a year. It has been rigorously tested and proven its worth on many occasions. So with a minor update today, and on the occasion of Kate Brooks's return to work, I decided to lift the version from 0.9.7 directly to 1.0.0. Congratulations TOAD! <ribbit><ribbit>
28 JAN 2010 / ind006: Various notes for the new year
10 OCT 2009 / ind006: Updated SiO 3mm/7mm maser databaseThe SiO 3mm maser database has been updated following recent reobservations of all masers.
17 JUN 2009 / edw126: ASAP point-and-click interfaceAn ASAP-GUI interface has been developed by Cormac Purcell and is available for download. It is also installed on kaputar. This GUI is useful for quick-look data reduction of position-switching data. It also has the capability to export the reduced data. To start the GUI type "python -i QuickData.py". Load in your data files with the "Browse" and "Load" buttons. Select the IF you wish to work on with "Current IF". First average in time the data for each of the two polarisations (2 blue boxes) and then average the two polarisations (2 green boxes).
06 APR 2009 / ind006: Updated SiO 3mm maser databaseThe SiO 3mm maser database has been updated following recent reobservations of all masers. These are the masers are used for pointing checks and updates of the Mopra telescope. The database now also enables the list to be sorted in order of angular separation from your favourite source.
06 APR 2009 / edw126: New version of ASAPVersion 2.3 of ASAP (the ATNF Spectral Analysis Package) was released on 13 March 2009. The new version allows output files to be written in a FITS format that can be read into the CLASS and XSpec software packages.
27 MAR 2009 / ind006: UPS and drive issuesDuring the last winter season, problems were experienced with the mains power supply and UPS unit at Mopra, which resulted in some observing time being lost. Over the summer, the UPS was replaced, and several additional UPS units have been added to spread the load across the site. Furthermore, the antenna drives, which were formerly powered via the UPS, have been taken off the UPS supply as there is some evidence that transient power spikes associated with the drives may have been a contributing cause to the UPS problems. As a result, any dropouts in mains supply will now have an immediate affect on the drives.
27 MAR 2009 / ind006: RFI spikes at 12mmIt has been reported there are two spikes, one at ~21.088 GHz that is ~1 channel (observing with 4096 channels across 137 MHz) wide, and another at 20.0528 GHz that is ~ 3 channels wide. Both spikes appear in both polarizations, neither move in frequency as the LO is changed.
We are aware of these and in the process of locating them and removing/replacing the culprit equipment. The spikes are not near any astrophysical lines.
Last update: Phil Edwards (03-Dec-2014) Original: Balt Indermuehle (20-March-2009)